You invested $700.00 compounded continously. After 20 years you had $2838.64. What was the interest rate?
Found 2 solutions by Nate, stanbon:
Answer by Nate(3500)

a(1 + r)^t
700(1 + r)^20 = 2,838.64
(1 + r)^20 = 4.0552
About 0.0725 or 7.25%

Answer by stanbon(48569)

You invested $700.00 compounded continously. After 20 years you had $2838.64. What was the interest rate?
--------------------
A(t)= Pe^(rt)
2838.64 = 700e^(20r)
e^(20r) = 4.0552
Take the natural log of both sides to get:
20r = ln(4.0552)
20r = 1.4
r = 0.07 or 7%
================
Cheers,
Stan H.